OpenStudy (anonymous):

Can someone help me with solving the logarithm? 1/2log(3)x=log(3)9

OpenStudy (anonymous):

can you rewrite it using the equation editor the wway you wrote it, it can be easily misunderstood and cannot not start with a wrong question.

OpenStudy (anonymous):

\[\frac{1}{2}\log_3(x)=\log_3(9)\]

OpenStudy (anonymous):

\[\log_3(9)=2\] since \(3^2=9\) so your equation is \[\frac{1}{2}\log_3(x)=2\]multiply by 2 and get \[\log_3(x)=4\] then rewrite in exponential form as \[3^4=x\]

OpenStudy (anonymous):

thank you